OpenStudy (christiannewsom):

Gina looks at the architectural plan of a four-walled room in which the walls meet each other at right angles. The length of one wall in the plan is 17 inches. The length of the diagonal of the floor of the room in the plan is approximately 18.79 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work.

OpenStudy (christiannewsom):

please help

OpenStudy (christiannewsom):

how do you find the answer i just need to know how to set it up and what to do

OpenStudy (christiannewsom):

ill medal and fan

OpenStudy (acannell):

The room is not a square. Using Pythagorean Theorem, a^2 + b^2 = c^2; we have the 'a' and the 'c'. If 'b' is the same as 'a', then it would be a square. 17^2 + b^2 = 18.79^2 289 + b^2 = 353 -289 -289 b^2 = 64 b = 8 It is not a square The dimensions are 8 x 17

OpenStudy (christiannewsom):

sorta thankyou

OpenStudy (acannell):

Sorry i keep confusing myself lol im making the question more hard than it actually is.

OpenStudy (christiannewsom):

lol its okay i understand it better because you said the Pythagorean theorem

OpenStudy (acannell):

Yea. Is it weird that i know this stuff and im in 7th grade???

OpenStudy (christiannewsom):

lol thanks

OpenStudy (acannell):

Your welcome!!! :) Hope i helped!!